The loudness level of a heavy snore is 69dB. The loudness level of a conversation is 60dB. The loudness level of a whisper is 30db.
a)How many times as lound as a conversation is a heavy snore?
b)How many times as lound as a whisper is a conversation?
I have to use this formula:L=10Log(I/Io)
Please include all steps. Thank you.

Answer by stanbon(57224)

The loudness level of a heavy snore is 69dB. The loudness level of a conversation is 60dB. The loudness level of a whisper is 30db.
a)How many times as lound as a conversation is a heavy snore?
b)How many times as lound as a whisper is a conversation
-----------
Snore: 10log(snore) = 69
log(snore) = 6.9
snore = 10^6.9
-------------
Whisper: 10log(whisper)= 30
log(whisper) = 3
whisper = 10^3
---------------------
Conversation: 10log(conversation)=60
log(conversation)= 6
conversation = 10^6
-----------------------
a)How many times as loud as a conversation is a heavy snore?
snore/conversation = 10^6.9/10^6 = 10^0.9 or 7.9 times as loud
-------------------------
b)How many times as loud as a whisper is a conversation
conversation/whisper = 10^6/10^3 = 10^3 or 1000 times as loud
---------------------------
Cheers,
Stan H.